Calculate the speed of sound in oxygen from the following data. The mass of 22.4 litre of oxygen at STP (T = 273 K and p = 1.0 × 105 N m-2) is 32 g, the molar heat capacity of oxygen at constant volume is Cv = 2.5 R and that at constant pressure is Cp = 3.5 R.

Grade:upto college level

1 Answers

Deepak Patra
9 years ago
Sol. Given that p = 1.0 × 10^5 N/m^2, T = 273 K, M = 32 g = 32 × 10^–3 kg V = 22.4 litre = 22.4 × 10^–3 m^3 C/C base v = r = 3.5 R / 2.5 R = 1.4 ⇒ V = √rp/f = √1.4 * 1.0 * 10^-5/32/22.4 = 310 m/s (because p = m/v)